Calculate the electric field at one corner off a square 80cm on a side if the other three corners are occupied by charges each of magnitude 18.2 x 10-4C.


1
Expert's answer
2021-05-23T18:36:23-0400

Gives

r=80cm 0.80m

Charge=18.2"\times20^{-4}C"

"E_1=\frac{Kq_1q_2}{r^2}"

"E_2=\frac{Kq_1q_2}{r^2}"

"E_3=\frac{Kq_1q_2}{r'^2}"

E1And E2 perpendiculer

Resultant

E'="\sqrt{E_1^2+E_2^2}"

"E'=\sqrt2E_1=\sqrt{2}E_2"

"E_3=\frac{kq_1q_2}{2r^2}"

Where r'="\sqrt{2}r"

"E_1=\frac{9\times10^9\times18.2\times10^{-4}\times18.2\times10^{-4}}{0.8^2}"

"E_1=46.59\times10^{3}N/C"

"E_1=E_2=46.59\times10^3N/C"


"E'=\sqrt2\times46.59\times10^3N/C=65.88\times10^3N/C"

"E_3=\frac{9\times10^9\times18.2\times10^{-4}\times18.2\times10^{-4}}{(\sqrt2\times0.8)^2}"

"E_3=23.295\times10^3N/C"

"E_{net}=E'+E_3"


"E_3=65.88\times10^3+23.295\times10^3=89.175\times10^3 N/C"